parallel to 2x+3y= -1 containing (4,5)
perpendicular to 2x+3y= -1 containing (3,2)

Answer by stanbon(75887)

parallel to 2x+3y= -1 containing (4,5)
y=(-2/3)x-(1/3)
Slope is -2/3
Any line parallel to the given line will have a slope of -2/3.
-------------------------
Point-Slope form:
y-5=(-2/3)(x-4)
y=(-2/3)x+8/4+20/4
y=(-2/3)x+7
-----------------------
perpendicular to 2x+3y= -1 containing (3,2)
y=(-2/3)x-1/3
The slope of the given line is -2/3
Any line perpendicular to the given line will have a slope of 3/2
----------------
Point-Slope form:
y-2=(3/2)(x-3)
y=(3/2)x-9/2+4/2
y=(3/2)x-(5/2)
-----------------
Cheers,
Stan H.